# Recording and Describing Poker Hands

### v3

@@ Abstract
This paper introduces the Poker Hand History (PHH) file format, designed to standardize the recording of poker hands across different game variants. Despite poker’s widespread popularity in the mainstream culture as a mind sport and its prominence in the field of artificial intelligence (AI) research as a benchmark for imperfect information AI agents, it lacks a consistent format that humans can use to document poker hands across different variants that can also easily be parsed by machines. To address this gap in the literature, we propose the PHH format which provides a concise human-readable machine-friendly representation of hand history that comprehensively captures various details of the hand, ranging from initial game parameters and actions to contextual parameters including but not limited to the venue, players, and time control information. In the supplementary, we provide 10,088 hands covering 11 different variants in the PHH format. The source code of the parser is available on GitHub: https://github.com/uoftcprg/pokerkit
@@ VI Required Fields
Forced bets blinds_or_straddles Array of integers or floats
@@
The showdown action is performed at the end of the hand or after all players go all-in. It represents showing of the hole cards to try to win the pot or after an all-in, or mucking of the hole cards. This action accepts an optional argument of cards which, if supplied, must all be known. The omission of this field represents mucking while the inclusion represents showing.
@@ VI-A Card Notation
A card in the PHH format can either be known or unknown. Each known card must be represented with two characters. The first character represents the rank of the card while the second character represents the suit, as tabulated in Tables V and VI. An unknown card must be represented with two question mark characters: “??”. Multiple cards must be represented by concatenating individual single-card representations without any separators or delimiters.
@@ VII Optional Fields
IANA time zone time_zone String
@@ IX Benchmarking
Note that the benchmarks of this sort for other hands included in the supplementary is impossible due to the fact that the Pluribus format only supports no-limit Texas hold ’em in specific initial configurations and no documentation is available on recording hands of different variants and configurations in PokerStars notation. In addition, the concepts used in some of the histories like big blind antes are only recently emerging and therefore not supported on the PokerStars platform.
@@
The performance of our open-source parser implementation, written in Python, is tabulated in Table IX. The hands in the supplementary were used to calculate the throughput and the parser process was run on a single CPU of Intel® Core™ i5-4690. All the file contents were loaded into memory prior to being parsed. We utilize Python’s built-in TOML file reader to parse the file and perform checks to validate the data. On average, it takes 0.147 milliseconds to parse a single hand. Note that no widely established open-source hand history parser for other file formats is available to facilitate a fair comparison.
@@ X Validation
Additionally, the parsed games for the Pluribus [5] hands were checked for consistency and correctness by comparing the final stacks with the payoffs provided in the dataset of Brown and Sandholm [5]. Similarly, the most of the televised poker hands from our historical selections and the final table of the 2019 WSOP Event #52: \$50,000 PPC on Day 5 had overlays of final player stacks in the source video from which we verified that the actions lead to the corresponding stack values. The actions were also checked for violations of the 2023 WSOP Tournament Rules [6].

### v4

@@ Abstract
This paper introduces the Poker Hand History (PHH) file format, designed to standardize the recording of poker hands across different game variants. Despite poker’s widespread popularity in the mainstream culture as a mind sport and its prominence in the field of artificial intelligence (AI) research as a benchmark for imperfect information AI agents, it lacks a consistent format that humans can use to document poker hands across different variants that can also easily be parsed by machines. To address this gap in the literature, we propose the PHH format which provides a concise human-readable machine-friendly representation of hand history that comprehensively captures various details of the hand, ranging from initial game parameters and actions to contextual parameters including but not limited to the venue, players, and time control information. In the supplementary, we provide 10,088 hands covering 11 different variants in the PHH format. The full specification is available on https://github.com/uoftcprg/phh-std
@@ VI Required Fields
Blinds/Straddles blinds_or_straddles Array of integers or floats
@@
The showdown action is performed at the end of the hand or after all players go all-in. It represents showing of the hole cards to try to win the pot or after an all-in, or mucking of the hole cards. This action accepts an optional argument of cards which, if supplied, must all be known. The omission represents mucking while the inclusion represents showing.
@@ VI-A Card Notation
A card in the PHH format can either be known or unknown. Each known card must be represented with two characters. The first character represents the rank of the card while the second character represents the suit, as tabulated in Tables V and VI. An unknown card must be represented with two question mark characters: “??”. Multiple cards must be represented by concatenating individual single-card representations without any separators or delimiters. For showdown actions only, if the hole cards being shown have already been specified during the corresponding player’s hole card dealing action, the cards can be written as a single dash character: “-”.
@@ VII Optional Fields
IANA time zone [18] time_zone String
@@ IX Benchmarking
Note that the benchmarks of this sort for other hands included in the supplementary is impossible due to the fact that the Pluribus format only supports no-limit Texas hold ’em in specific initial configurations and no documentation is available on recording hands of different variants and configurations in PokerStars notation. In addition, the concepts used in some of the histories like big blind antes are only recently emerging and therefore not supported on PokerStars.
@@
The performance of our open-source parser implementation in PokerKit [9] is tabulated in Table IX. The hands in the supplementary were used to calculate the throughput and the parser process was run on a single CPU of Intel® Core™ i5-4690. All the file contents were loaded into memory prior to being parsed. We utilize Python’s built-in TOML file reader to parse the file and perform checks to validate the data. On average, it takes 0.147 milliseconds to parse a single hand. Note that no widely established open-source hand history parser for other file formats is available to compare with.
@@ X Validation
Additionally, the parsed games for the Pluribus [5] hands were checked for consistency and correctness by comparing the final stacks with the payoffs provided in the dataset of Brown and Sandholm [5]. Similarly, the most of the televised poker hands from our historical selections and the final table of the 2023 WSOP Event #43: \$50,000 PPC on Day 5 [12] had overlays of final player stacks in the source video from which we verified that the actions lead to the corresponding stack values. The actions were also checked for violations of the 2023 WSOP Tournament Rules [6].